40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6611     Entry time: Mon May 7 01:07:58 2012
Author: Den 
Type: Update 
Category: IOO 
Subject: seismic in mcf 

I tried to figure out where additional (to seismic) noise enters to MC_F, so the coherence below 1 Hz is low (~0.2-0.5). I've examined noise in the path

PD -> DEMOD -> MC BOARD -> FSS -> PZT Controller -> LASER

  • I terminated ADC and measured its noise
  • connected MC BOARD OUT to ADC, terminated INPUT1 of the MC BOARD and measured noise
  • connected DEMOD Q OUT to MC BOARD INPUT1, terminated PD INPUT on the DEMODULATOR and measured noise
  • connected PD to DEMOD, blocked the beam incident to the PD and measured noise

pd-mcf.jpeg

 MC BOARD noise shows up only below 0.1 Hz and at 10 Hz where the whitening filter starts to work. SNR is ~2 at 4 Hz, so we might want to slightly improve whitening filter. But other then that path PD->DEMOD->MC BOARD is not responsible for additional noises below 1 Hz.

Next I connected SR785 to the laser and measured the closed loop feedback signal while MC was locked.

fb.png     coh_err_fb.png

Coherence between MC BOARD OUT1 and feedback signal is high enough to assume that FSS and PZT controller are also not responsible for additional noise.

From the other hand MC2 SUSPOS measured by OSEMS shows good coherence with GUR1_X

mc2_suspos.jpg

That means that MC2 is indeed driven by seismic motion. In order to figure out if this is the case for MC1 and MC3, I rotated GUR2 by 45 degrees. When it will calm down, I'll measure coherence between OSEMs and seismic motion.

ELOG V3.1.3-